OpenStudy (anonymous):

find the general solution of the differential equation and use it to determine solutions as t -> infinity y'+3y=t+e^(-2t)

OpenStudy (amistre64):

what is the homogenous solution?

OpenStudy (amistre64):

y' + 3y = 0

OpenStudy (anonymous):

\[e ^{3t}y=\int\limits_{?}^{?}e ^{3t}*(t+e ^{-2t})dt\] is as far as i got

OpenStudy (anonymous):

my integrating factor was \[\mu(t)= e ^{\int\limits(3(t)dt}\]

OpenStudy (amistre64):

\[y_h=A~e^{rt}\]\[r+3=0~:~r=-3\]\[y_h=A~e^{-3t}\] \[y_p=A(t)~e^{-3t}\]\[y'_p=A'(t)~e^{-3t}-3A(t)e^{-3t}~:~A'(t)~e^{-3t}=0\]\[y'_p=-3A(t)e^{-3t}\] its been awhile since i tried it using an I.F.

OpenStudy (anonymous):

well the answer is \[y=ce ^{-3t}+\frac{ t }{ 3 }-\frac{ 1 }{ 9 }+e ^{-2t}\]

OpenStudy (anonymous):

i got it from the back of the book but have no idea how to get there from where i am now

OpenStudy (amistre64):

\[A'(t)e^{-3t}=t+e^{-2t}\] \[A'(t)=te^{3t}+e^{(-2+3)t}\] \[\int A'(t)dt=\int te^{3t}+e^{(-2+3)t}dt\] \[\int A'(t)dt=\frac{t}{3}e^3-\frac{1}{9}e^{3t}+e^{t}+C\] \[y=c~e^{-3t}+\frac{t}{3}e^3-\frac{1}{9}e^{3t}+e^{t}+C\] pfft, i can get someplace close to it, but i doubt my method was sound

OpenStudy (anonymous):

im still not understanding where u got this A from

OpenStudy (amistre64):

\[y=c~e^{-3t}+(\frac{t}{3}-\frac{1}{9}+e^{-2t})e^{-3t}\]

OpenStudy (amistre64):

A is just something i used to call "c" by

OpenStudy (anonymous):

o ok thank you very much

OpenStudy (amistre64):

somehow i either get an extra \(e^{-3t}\), but it might be appropriate ... i just dont see how lol

OpenStudy (anonymous):

was my integrating factor and my first equation setup before i integrated right?

OpenStudy (amistre64):

should used e^(-3t), but from what i can recall, it looks like a good setup otherwise

OpenStudy (anonymous):

alright but i still dont understand how you do this: \[\int\limits(e ^{3t}(t+e ^{-2t})dt\]

OpenStudy (amistre64):

y = Ce^(-3t) y' = -3Ce^(-3t) y' + 3y = 0 -3Ce^(-3t) + 3Ce^(-3t) = 0

OpenStudy (amistre64):

had connection issues with the site

OpenStudy (amistre64):

distribute it thru and integrate the pieces

OpenStudy (anonymous):

i got \[e ^{3t}t+e ^{t}\]

OpenStudy (amistre64):

good, not the last term is simple enough, the first term takes it by parts

OpenStudy (amistre64):

*now

OpenStudy (amistre64):

u = t v = e^3t/3 du = dt dv = e^3t dt {S} te^(3t) dt = t e^3t/3 - {S} e^3t/3 dt {S} te^(3t) dt = t e^3t/3 - e^3t/9 {S} te^(3t) dt = e^3t( t/3 - 1/9)

OpenStudy (amistre64):

i spose +C is in order :) e^3t( t/3 - 1/9) + e^t + C now divide off that e^3t from the front ( t/3 - 1/9) + e^-2t + Ce^-3t

OpenStudy (anonymous):

thanks for your help but im totally lost, i think ill just go back to clac 2

OpenStudy (amistre64):

what part did you get lost at? or would this look better in the coding?

OpenStudy (anonymous):

well i got te^3t+ e^t + C and u got e^3t (t/3-1/9) + C

OpenStudy (amistre64):

\[e^{3t}y=\int te^{3t}dt+\int e^tdt\] \[e^{3t}y=\int te^{3t}dt+e^t+c\] \[e^{3t}y=\frac{1}{3}te^{3t}-\frac{1}{3}\int e^{3t}dt+e^t+c\] \[e^{3t}y=\frac{1}{3}te^{3t}-\frac{1}{3}\frac{1}{3}e^{3t}+c+e^t+c\] \[e^{3t}y=\frac{t}{3}e^{3t}-\frac{1}{9}e^{3t}+e^t+C\]divide off that front e^3t\[y=\frac{t}{3}\frac{e^{3t}}{e^{3t}}-\frac{1}{9}\frac{e^{3t}}{e^{3t}}+\frac{e^t}{e^{3t}}+\frac{C}{e^{3t}}\] \[y=\frac{t}{3}-\frac{1}{9}+e^{-2t}+Ce^{-3t}\]

OpenStudy (anonymous):

I got what "the back of the book" got, but I don't understand the last part of the question: \(\quad\)"Determine solutions as \(t\rightarrow\infty\)" The expression \(y=\frac{1}{3}t-\frac{1}{9}+e^{-2t}+ke^{-3t}\) clearly diverges as \(t\rightarrow\infty\)

OpenStudy (anonymous):

omg yes finally thank you soooooo much amistre64, i may have hope after all

OpenStudy (amistre64):

:) good luck